# Maximum entropy models of neuronal populations at and off criticality

### v1

@@ I Introduction
ME modeling has proven a powerful approach to study the dynamics of biological neural networks [schneidman_weak_2006, tkacik_spin_2009, tkacik_searching_2014]. Spiking activity of neurons can be described as a binary process, $\sigma(t)$, in which $\sigma=\pm 1$ represents the state of the neuron at a given time $t$, i.e. spiking for $\sigma=1$ and silent for $\sigma=-1$ [rieke_spikes_1999]. This approach allows one to define an empirical distribution of binary activity patterns, $P(\bf{\sigma})$, which specifies the probability of observing a given spiking pattern in a population of neurons. For NN binary neurons, the distribution $P(\bf{\sigma})$ for all $2^{N}$ possible spiking patterns (or states) fully characterizes population dynamics. However, estimating $P(\bf{\sigma})$ directly from data is often impractical since the number of possible states grows exponentially with the number of neurons [mora_are_2011]. Maximum entropy modeling offers a possible solution to this problem by constructing the least-biased probability distribution that matches selected statistics, such as individual firing rates and pairwise correlations, while maximizing the entropy of $P(\bf{\sigma})$. The resulting distribution is mathematically equivalent to the Boltzmann distribution [schneidman_weak_2006, tkacik_spin_2009, gardella_modeling_2019], which can then be analyzed using thermodynamic tools, such as the fluctuation-dissipation theorem [tkacik_thermodynamics_2015, sampaio_filho_ising-like_2024, simoes_thermodynamic_2024].
@@ II.1 Integrate-and-fire neuronal network model
We consider an integrate-and-fire (IF) model on a scale-free, directed network with short- and long-term plasticity, and refractory period of one timestep during which neurons remain inactive after firing [michiels_van_kessenich_critical_2018]. We implement networks with different number of neurons, NN, placed within a cube of side $L$, but keep the density $N/L^{3}=0.016$ constant [nandi_scaling_2022]. A fraction $p_{\textrm{in}}=20\%$ of neurons is inhibitory [zeraati_self-organization_2021]. The out-degree $k$ of each neuron follows a power-law distribution, $P(k)\propto k^{-2}$, with $k\in[2,20]$ for systems with $N\leq 100$ and $k\in[2,100]$ otherwise. The connection probability between two neurons decays exponentially with the Euclidean distance $r$, $P(r)\propto e^{-r/r_{0}}$, where $r_{0}=5$ [roerig_relationships_2002]. The resting potential of each neuron $i$ is set at $v_{i}=0$. A neuron fires when $v_{i}\geq v_{c}=1$, transmitting signals to all its post-synaptic neurons $j$ according to the following equation:
@@
where ++ and - are for excitatory and inhibitory pre-synaptic neurons, $g_{ij}$ is the strength of the synapsis connecting $i$ to $j$, and $u_{i}$ indicates the synaptic resources of the pre-synaptic neuron ii. The constant $\delta u=0.05$ controls the fraction of neurotransmitters released [ikeda_counting_2009]. Here, the synaptic resources $u_{i}(t)\in[0,1]$ and the synaptic strengths $g_{ij}\in[10^{-5},1]$ change in time to model the short- and long-term plasticity, respectively.
@@
After each avalanche, the pool of neurotransmitters of each neuron, $u_{i}$, is recovered by an amount $\delta u_{\textrm{rec}}$, i.e. $u_{i}(t)\rightarrow u_{i}(t)+\delta u_{\textrm{rec}}$. Recovery is implemented between avalanches according to the separation of time scales in self-organized models, which considers avalanches as almost instantaneous events. For a given NN, the system can be set to criticality by tuning $\delta u_{\textrm{rec}}$ to a certain value $\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$. At criticality, the system exhibits avalanches with size $S$ and duration $D$ distributed according to power-laws whose cut-off scales with the system size NN. Setting $\delta u_{\textrm{rec}}<\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ leads to subcritical dynamics, characterized by an exponential decay in the distributions of avalanche sizes, $P(S)$, and avalanche durations, $P(D)$. Conversely, for $\delta u_{\textrm{rec}}>\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ the system is in a super-critical state where there is a sharp increase in large and long avalanches, leading to the appearance of local maxima in $P(D)$ and in $P(S)$ around the power-law cut-off (see supplementary Fig. S1) for the distributions and Table S1 for the set of values $\delta u_{\textrm{rec}}^{\textrm{crit}}$). For simulations of subcritical and supercritical IF networks, we set $\delta u_{\textrm{rec}}(N)=0.1\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ and $\delta u_{\textrm{rec}}(N)=10\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$, respectively.
@@
For each system size NN and state of the dynamics (subcritical, critical, and supercritical), we generate five independent time series of neuronal spikes, each corresponding to a different realization of the network configuration. We refer the reader to the SI for a comprehensive account of the avalanche size and duration distributions in each network state (Fig. S1).
@@ II.3 Time-averaged firing statistics
where $\delta_{K,K^{\prime}(k)}$ is the Kronecker delta function and $K^{\prime}(k)=\sum_{i=1}^{N}\left(\sigma_{i}(k)+1\right)/2$ counts the number of neurons or electrodes that fired during time bin $k$. Notice that, since $P(K)$ is a distribution, only NN out of the $N+1$ values of $P(K)$ are independent because of the normalization condition $\sum_{K=0}^{N}P(K)=1$.
@@ II.3.1 Allometry of firing rates
The firing rates $r_{i}=(\langle\sigma_{i}\rangle+1)/2\Delta t_{b}$ define the total firing rate $\langle n_{a}\rangle=\sum_{i=1}^{N}r_{i}$, an interesting quantity that exhibits allometric scaling—a non-trivial scaling relationship—with the number of neurons, $\langle n_{a}\rangle\propto N^{\eta}$, with an allometric exponent $\eta<1$ [simoes_allometric_2025, karbowski_thermodynamic_2009], which can be derived from the finite-size scaling of neuronal avalanches [simoes_allometric_2025]. In the IF model, even if the network operates off criticality, $\langle n_{a}\rangle$ scales sublinearly with NN (SI, Fig. S3), with a slightly lower or higher allometric exponent if the dynamics is subcritical or supercritical, respectively. The main difference between the curves is the absolute value $\langle n_{a}\rangle$, larger in the supercritical networks and smaller in subcritical ones (SI, Fig. S3). Provided that the average quiet time between successive avalanches does not diverge faster than the average avalanche duration in the limit $N\rightarrow\infty$ (SI, Fig. S4), the robustness of the allometric exponent across the states of the dynamics could be explained by the fact that avalanche size and duration distributions always have an intermediate power-law regime with similar exponents in the IF model (see supplementary Fig. S5), with a cutoff that scales similarly with NN, independently of the tuning parameter $\delta u_{\textrm{rec}}$. However, it is important to note that in the subcritical regime the distributions are nevertheless closer to exponential-like, decaying well before reaching system-size avalanches, while in the supercritical regime the distributions show an increased probability of large, system-spanning avalanches, as evidenced by a local maximum near the system size.
@@ III Maximum Entropy Modeling
where $n$ is the iteration index and $\theta(n)\propto n^{-\alpha}$ is a decreasing learning rate, with the value of $\alpha>0$ that can be adjusted depending on NN and the state of the dynamics of each dataset (see Table S2 in the SI).
@@ III.1 $K$-pairwise ME models
We are interested in finding the ME distribution $P_{\textrm{MEM}}(\bm{\sigma})$ that constrains the NN average local activities $\langle\sigma_{i}\rangle$, the $N\cdot(N-1)/2$ pairwise correlations $\langle\sigma_{i}\sigma_{j}\rangle$, which in turn constrain the correlation functions $C_{ij}$, and the synchronous firing probability distribution $P(K)$. This corresponds to NN Lagrange multipliers for $f_{i}(\bm{\sigma})=\sigma_{i}$, denoted by $h_{i}$, $N\cdot(N-1)/2$ for $f_{ij}(\bm{\sigma})=\sigma_{i}\sigma_{j}$, denoted by $J_{ij}$, and $N+1$ for $f_{K}(\bm{\sigma})=\delta_{K,K^{\prime}(\bm{\sigma})}$, denoted by $V_{K}$. As mentioned previously, only NN out of the $N+1$ possible values of $P(K)$ are independent, so one only needs to fit at most NN parameters $V_{K}$. Therefore, we set $V_{K=0}$ to zero. The Hamiltonian in (9) then reads,
@@
We start with $h_{i}(n=1)=\langle\sigma_{i}\rangle_{\textrm{data}}$, $J_{ij}(n=1)=0$ and $V_{K}(n=1)=0$ and then iterate equations (15)-(17) typically until $n\sim 2\cdot 10^{5}$. At each iteration, quantities are averaged over $M_{c}=3\cdot 10^{5}$ spin configurations using the Metropolis algorithm. We disregard the first $150N$ Monte Carlo iterations in order to reduce correlations with the initial state. To avoid divergence issues due to poor sampling at small $P(K)$, we only fit the $V_{K}$ associated with $P(K)>10^{-4}$ for experimental data or $P(K)>10^{-5}$ for the numerical data, and set all other $V_{K}=0$. We use a smaller threshold for numerical data because $P(K)$ is better sampled thanks to the larger number of time bins that can be considered—experimental data instead have a limited $\approx 1$h duration. At the end of the learning routine, we study the inferred Ising-like models with the set of fitted parameters $\{h_{i},J_{ij},V_{K}\}$ by averaging over an increased amount of spin configurations $M_{c}=3\cdot 10^{6}$, averaged over 100100 random initial spin configurations, to reduce error bars.
@@ III.3 Predictive capability of the K-pairwise Ising models
Figure 4: Predictive capability of the K-pairwise Ising models inferred from numerical and experimental data. Comparison of the three-point correlation functions $T_{ijk}$ between the data (x-axes) and the prediction of the maximum entropy distribution (Eq. (9)) (y-axes), for IF networks with $N=100$ neurons in the subcritical (a), critical (b), and supercritical (c) states, as well as experimental recordings from cultures treated with a combination of AP5/DNQX (d), cultures in baseline, no-drug condition (e) and cultures treated with PTX (f). The blue dashed lines indicate the bisector $y=x$. For the original data (x-axes), results are averaged over $N_{b}=10^{7}$ time bins for the IF model, and over $N_{b}\approx 1.4\cdot 10^{5}$ time bins for the experimental datasets. For the sampling of the distribution $P_{\textrm{MEM}}(\bm{\sigma})$ from Eq. (9) (y-axes), using the respective learned parameters shown in Fig. 2, results are averaged over $M_{c}=3\cdot 10^{6}$ spin configurations, and over 100100 random initial spin configurations. Error bars represent the standard error of the mean, and, with the exception of the subcritical cases (a) and (d), are smaller than or equal to the symbol size.
@@ III.4 Thermodynamics of the K-pairwise Ising models
Figure 6: Thermodynamic response functions of the K-pairwise Ising-like models inferred from experimental data. Same as in Fig. 5 but for data recorded in cortex slice cultures ($N=60$ electrodes) treated with a combination of AP5/DNQX (left column), in baseline, no-drug cultures (center column) and in cultures treated with PTX (right column) The gray background indicates the temperature regime where Metropolis Monte Carlo results depend on the initial spin configuration. The colored shaded areas around the curves of $C_{v}$ and $\chi$ is the standard error obtained from 55 experimental subsamples of the same cortical culture.
@@
Given a set of NN spins, the total magnetization $M(\bm{\sigma})=\sum_{i=1}^{N}\sigma_{i}$ and the energy $H(\bm{\sigma})$ defined in Eq. (14) can be obtained for different temperatures. From the fluctuations of these quantities, according to the fluctuation-dissipation theorem, the isothermal magnetic susceptibility $\chi$ and heat capacity $C_{v}$ can be calculated as functions of $T$,
@@
The K-pairwise Ising models inferred from real and IF networks in the critical state present maxima of the heat capacity and susceptibility at $T=T_{\textrm{max}}$ slightly larger than one (see SI, Table S3 for the values of $T_{\textrm{max}}$). In the models inferred from critical IF networks, these maxima increase faster-than-linearly with the system size NN for the critical case (Figs. 5b,e), scaling as $\textrm{max}[C_{v}]\propto N^{a}$ with $a=1.41\pm 0.06$ and $\textrm{max}[\chi]\propto N^{b}$ with $b=1.57\pm 0.09$, where the exponents were estimated from a least-square fit of the data and their errors calculated by bootstrapping. Importantly, for the critical IF model, $T_{\textrm{max}}$ consistently approaches $T=1$ as NN increases, suggesting that in the thermodynamic limit $N\to\infty$ the inferred Ising models operate at a critical point. This is consistent with the original criticality of the IF model data, and in line with results from pairwise Ising models that were inferred by constraining only $\{\langle\sigma_{i}\rangle_{\textrm{data}}\}$ and $\{C_{ij}^{\textrm{data}}\}$ [simoes_thermodynamic_2024].
Notably, the models inferred from supercritical IF (Fig. 5c,f) and real neuronal networks (Fig. 6c,f) also display maxima in $C_{v}$ and $\chi$ close to $T=1$. In this case, the maximum of $C_{v}$ increases with the number of IF neurons, NN, faster than in the critical case, i.e. $\textrm{max}[C_{v}]\propto N^{a}$ with $a=1.70\pm 0.06$. Unlike $C_{v}$, the maximum of $\chi$ follows a scaling consistent with the critical case, i.e. $\textrm{max}[\chi]\propto N^{b}$ with $b=1.77\pm 0.12$.
@@
To conclude the analysis of the thermodynamics of the inferred Ising models, we examine the low temperature behavior of $C_{v}$ and $\chi$. For $T<T^{*}<1$, the numerical values of $C_{v}$ and $\chi$ depend on the initial spin configuration (see SI, Figs. S13 and S14). This interval $T<T^{*}$ is indicated by a gray-shaded background in Figs. 5 and 6 for the largest system sizes considered ($N=100$ neurons in the IF model and $N=60$ electrodes for experimental data). This behavior suggests the presence of a complex phase below the phase transition $T<T_{\textrm{max}}$, analogous to the spin-glass phase in the Sherrington-Kirkpatrick model [sherrington_solvable_1975], where many frustrated spin configurations are present and thermal fluctuations are insufficient to drive the system from a random initial state to the ground state within feasible Monte Carlo simulation times [newman_monte_1999, olsen_quality_2024]. The divergence of $C_{v}$ and $\chi$ as $T\to 0$ is therefore a numerical artifact, likely the result of averaging over different metastable states with varying energies $H(\bm{\sigma})$ and magnetizations $M(\bm{\sigma})$, resulting in large values for the fluctuations in Eqs. (20) and (21). The temperature $T^{*}$ depends on the system size, increasing with NN, and on the state of the dynamics of the IF networks, being smallest for subcritical IF networks and largest for supercritical ones (see SI, Fig. S13 and Table S4). For low temperature $T\rightarrow 0$, Ising-like models with disorder in the parameters, such as the Sherrington–Kirkpatrick spin glass model, exhibit metastable states separated by free-energy barriers that scale with NN [aspelmeier_free-energy_2006]. If K‐pairwise models fitted to neuronal data would also exhibit similar free‐energy barriers that scale with NN at low temperature, then the shift of $T^{*}$ to higher temperatures with system size could be explained by the increasing thermal energy needed to overcome these barriers as NN increases.

### v2

@@ I Introduction
ME modeling has proven a powerful approach to study the dynamics of biological neural networks [schneidman_weak_2006, tkacik_spin_2009, tkacik_searching_2014]. Spiking activity of neurons can be described as a binary process, $\sigma(t)$, in which $\sigma=\pm 1$ represents the state of the neuron at a given time $t$, i.e. spiking for $\sigma=1$ and silent for $\sigma=-1$ [rieke_spikes_1999]. This approach allows one to define an empirical distribution of binary activity patterns, $P(\bf{\sigma})$, which specifies the probability of observing a given spiking pattern in a population of neurons. For $N$ binary neurons, the distribution $P(\bf{\sigma})$ for all $2^{N}$ possible spiking patterns (or states) fully characterizes population dynamics. However, estimating $P(\bf{\sigma})$ directly from data is often impractical since the number of possible states grows exponentially with the number of neurons [mora_are_2011]. Maximum entropy modeling offers a possible solution to this problem by constructing the least-biased probability distribution that matches selected statistics, such as individual firing rates and pairwise correlations, while maximizing the entropy of $P(\bf{\sigma})$. The resulting distribution is mathematically equivalent to the Boltzmann distribution [schneidman_weak_2006, tkacik_spin_2009, gardella_modeling_2019], which can then be analyzed using thermodynamic tools, such as the fluctuation-dissipation theorem [tkacik_thermodynamics_2015, sampaio_filho_ising-like_2024, simoes_thermodynamic_2024].
@@ II.1 Integrate-and-fire neuronal network model
We consider an integrate-and-fire (IF) model on a scale-free, directed network with short- and long-term plasticity, and refractory period of one timestep during which neurons remain inactive after firing [michiels_van_kessenich_critical_2018]. We implement networks with different number of neurons, $N$, placed within a cube of side $L$, but keep the density $N/L^{3}=0.016$ constant [nandi_scaling_2022]. A fraction $p_{\textrm{in}}=20\%$ of neurons is inhibitory [zeraati_self-organization_2021]. The out-degree $k$ of each neuron follows a power-law distribution, $P(k)\propto k^{-2}$, with $k\in[2,20]$ for systems with $N\leq 100$ and $k\in[2,100]$ otherwise. The connection probability between two neurons decays exponentially with the Euclidean distance $r$, $P(r)\propto e^{-r/r_{0}}$, where $r_{0}=5$ [roerig_relationships_2002]. The resting potential of each neuron $i$ is set at $v_{i}=0$. A neuron fires when $v_{i}\geq v_{c}=1$, transmitting signals to all its post-synaptic neurons $j$ according to the following equation:
@@
where ++ and - are for excitatory and inhibitory pre-synaptic neurons, $g_{ij}$ is the strength of the synapsis connecting $i$ to $j$, and $u_{i}$ indicates the synaptic resources of the pre-synaptic neuron $i$. The constant $\delta u=0.05$ controls the fraction of neurotransmitters released [ikeda_counting_2009]. Here, the synaptic resources $u_{i}(t)\in[0,1]$ and the synaptic strengths $g_{ij}\in[10^{-5},1]$ change in time to model the short- and long-term plasticity, respectively.
@@
After each avalanche, the pool of neurotransmitters of each neuron, $u_{i}$, is recovered by an amount $\delta u_{\textrm{rec}}$, i.e. $u_{i}(t)\rightarrow u_{i}(t)+\delta u_{\textrm{rec}}$. Recovery is implemented between avalanches according to the separation of time scales in self-organized models, which considers avalanches as almost instantaneous events. For a given $N$, the system can be set to criticality by tuning $\delta u_{\textrm{rec}}$ to a certain value $\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$. At criticality, the system exhibits avalanches with size $S$ and duration $D$ distributed according to power-laws whose cut-off scales with the system size $N$. Setting $\delta u_{\textrm{rec}}<\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ leads to subcritical dynamics, characterized by an exponential decay in the distributions of avalanche sizes, $P(S)$, and avalanche durations, $P(D)$. Conversely, for $\delta u_{\textrm{rec}}>\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ the system is in a super-critical state where there is a sharp increase in large and long avalanches, leading to the appearance of local maxima in $P(D)$ and in $P(S)$ around the power-law cut-off (see supplementary Fig. S1) for the distributions and Table S1 for the set of values $\delta u_{\textrm{rec}}^{\textrm{crit}}$). For simulations of subcritical and supercritical IF networks, we set $\delta u_{\textrm{rec}}(N)=0.1\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$ and $\delta u_{\textrm{rec}}(N)=10\delta u_{\textrm{rec}}^{\textrm{crit}}(N)$, respectively.
@@
For each system size $N$ and state of the dynamics (subcritical, critical, and supercritical), we generate five independent time series of neuronal spikes, each corresponding to a different realization of the network configuration. We refer the reader to the SI for a comprehensive account of the avalanche size and duration distributions in each network state (Fig. S1).
@@ II.3 Time-averaged firing statistics
where $\delta_{K,K^{\prime}(k)}$ is the Kronecker delta function and $K^{\prime}(k)=\sum_{i=1}^{N}\left(\sigma_{i}(k)+1\right)/2$ counts the number of neurons or electrodes that fired during time bin $k$. Notice that, since $P(K)$ is a distribution, only $N$ out of the $N+1$ values of $P(K)$ are independent because of the normalization condition $\sum_{K=0}^{N}P(K)=1$.
@@ II.3.1 Allometry of firing rates
The firing rates $r_{i}=(\langle\sigma_{i}\rangle+1)/2\Delta t_{b}$ define the total firing rate $\langle n_{a}\rangle=\sum_{i=1}^{N}r_{i}$, an interesting quantity that exhibits allometric scaling—a non-trivial scaling relationship—with the number of neurons, $\langle n_{a}\rangle\propto N^{\eta}$, with an allometric exponent $\eta<1$ [simoes_allometric_2025, karbowski_thermodynamic_2009], which can be derived from the finite-size scaling of neuronal avalanches [simoes_allometric_2025]. In the IF model, even if the network operates off criticality, $\langle n_{a}\rangle$ scales sublinearly with $N$ (SI, Fig. S3), with a slightly lower or higher allometric exponent if the dynamics is subcritical or supercritical, respectively. The main difference between the curves is the absolute value $\langle n_{a}\rangle$, larger in the supercritical networks and smaller in subcritical ones (SI, Fig. S3). Provided that the average quiet time between successive avalanches does not diverge faster than the average avalanche duration in the limit $N\rightarrow\infty$ (SI, Fig. S4), the robustness of the allometric exponent across the states of the dynamics could be explained by the fact that avalanche size and duration distributions always have an intermediate power-law regime with similar exponents in the IF model (see supplementary Fig. S5), with a cutoff that scales similarly with $N$, independently of the tuning parameter $\delta u_{\textrm{rec}}$. However, it is important to note that in the subcritical regime the distributions are nevertheless closer to exponential-like, decaying well before reaching system-size avalanches, while in the supercritical regime the distributions show an increased probability of large, system-spanning avalanches, as evidenced by a local maximum near the system size.
@@ III Maximum Entropy Modeling
where $n$ is the iteration index and $\theta(n)\propto n^{-\alpha}$ is a decreasing learning rate, with the value of $\alpha>0$ that can be adjusted depending on $N$ and the state of the dynamics of each dataset (see Table S2 in the SI).
@@ III.1 $K$-pairwise ME models
We are interested in finding the ME distribution $P_{\textrm{MEM}}(\bm{\sigma})$ that constrains the $N$ average local activities $\langle\sigma_{i}\rangle$, the $N\cdot(N-1)/2$ pairwise correlations $\langle\sigma_{i}\sigma_{j}\rangle$, which in turn constrain the correlation functions $C_{ij}$, and the synchronous firing probability distribution $P(K)$. This corresponds to $N$ Lagrange multipliers for $f_{i}(\bm{\sigma})=\sigma_{i}$, denoted by $h_{i}$, $N\cdot(N-1)/2$ for $f_{ij}(\bm{\sigma})=\sigma_{i}\sigma_{j}$, denoted by $J_{ij}$, and $N+1$ for $f_{K}(\bm{\sigma})=\delta_{K,K^{\prime}(\bm{\sigma})}$, denoted by $V_{K}$. As mentioned previously, only $N$ out of the $N+1$ possible values of $P(K)$ are independent, so one only needs to fit at most $N$ parameters $V_{K}$. Therefore, we set $V_{K=0}$ to zero. The Hamiltonian in (9) then reads,
@@
We start with $h_{i}(n=1)=\langle\sigma_{i}\rangle_{\textrm{data}}$, $J_{ij}(n=1)=0$ and $V_{K}(n=1)=0$ and then iterate equations (15)-(17) typically until $n\sim 2\cdot 10^{5}$. At each iteration, quantities are averaged over $M_{c}=3\cdot 10^{5}$ spin configurations using the Metropolis algorithm. We disregard the first $150N$ Monte Carlo iterations in order to reduce correlations with the initial state. To avoid divergence issues due to poor sampling at small $P(K)$, we only fit the $V_{K}$ associated with $P(K)>10^{-4}$ for experimental data or $P(K)>10^{-5}$ for the numerical data, and set all other $V_{K}=0$. We use a smaller threshold for numerical data because $P(K)$ is better sampled thanks to the larger number of time bins that can be considered—experimental data instead have a limited $\approx 1$h duration. At the end of the learning routine, we study the inferred Ising-like models with the set of fitted parameters $\{h_{i},J_{ij},V_{K}\}$ by averaging over an increased amount of spin configurations $M_{c}=3\cdot 10^{6}$, averaged over $100$ random initial spin configurations, to reduce error bars.
@@ III.3 Predictive capability of the K-pairwise Ising models
Figure 4: Predictive capability of the K-pairwise Ising models inferred from numerical and experimental data. Comparison of the three-point correlation functions $T_{ijk}$ between the data (x-axes) and the prediction of the maximum entropy distribution (Eq. (9)) (y-axes), for IF networks with $N=100$ neurons in the subcritical (a), critical (b), and supercritical (c) states, as well as experimental recordings from cultures treated with a combination of AP5/DNQX (d), cultures in baseline, no-drug condition (e) and cultures treated with PTX (f). The blue dashed lines indicate the bisector $y=x$. For the original data (x-axes), results are averaged over $N_{b}=10^{7}$ time bins for the IF model, and over $N_{b}\approx 1.4\cdot 10^{5}$ time bins for the experimental datasets. For the sampling of the distribution $P_{\textrm{MEM}}(\bm{\sigma})$ from Eq. (9) (y-axes), using the respective learned parameters shown in Fig. 2, results are averaged over $M_{c}=3\cdot 10^{6}$ spin configurations, and over $100$ random initial spin configurations. Error bars represent the standard error of the mean, and, with the exception of the subcritical cases (a) and (d), are smaller than or equal to the symbol size.
@@ III.4 Thermodynamics of the K-pairwise Ising models
Figure 6: Thermodynamic response functions of the K-pairwise Ising-like models inferred from experimental data. Same as in Fig. 5 but for data recorded in cortex slice cultures ($N=60$ electrodes) treated with a combination of AP5/DNQX (left column), in baseline, no-drug cultures (center column) and in cultures treated with PTX (right column) The gray background indicates the temperature regime where Metropolis Monte Carlo results depend on the initial spin configuration. The colored shaded areas around the curves of $C_{v}$ and $\chi$ is the standard error obtained from $5$ experimental subsamples of the same cortical culture.
@@
Given a set of $N$ spins, the total magnetization $M(\bm{\sigma})=\sum_{i=1}^{N}\sigma_{i}$ and the energy $H(\bm{\sigma})$ defined in Eq. (14) can be obtained for different temperatures. From the fluctuations of these quantities, according to the fluctuation-dissipation theorem, the isothermal magnetic susceptibility $\chi$ and heat capacity $C_{v}$ can be calculated as functions of $T$,
@@
The K-pairwise Ising models inferred from real and IF networks in the critical state present maxima of the heat capacity and susceptibility at $T=T_{\textrm{max}}$ slightly larger than one (see SI, Table S3 for the values of $T_{\textrm{max}}$). In the models inferred from critical IF networks, these maxima increase faster-than-linearly with the system size $N$ for the critical case (Figs. 5b,e), scaling as $\textrm{max}[C_{v}]\propto N^{a}$ with $a=1.41\pm 0.06$ and $\textrm{max}[\chi]\propto N^{b}$ with $b=1.57\pm 0.09$, where the exponents were estimated from a least-square fit of the data and their errors calculated by bootstrapping. Importantly, for the critical IF model, $T_{\textrm{max}}$ consistently approaches $T=1$ as $N$ increases, suggesting that in the thermodynamic limit $N\to\infty$ the inferred Ising models operate at a critical point. This is consistent with the original criticality of the IF model data, and in line with results from pairwise Ising models that were inferred by constraining only $\{\langle\sigma_{i}\rangle_{\textrm{data}}\}$ and $\{C_{ij}^{\textrm{data}}\}$ [simoes_thermodynamic_2024].
Notably, the models inferred from supercritical IF (Fig. 5c,f) and real neuronal networks (Fig. 6c,f) also display maxima in $C_{v}$ and $\chi$ close to $T=1$. In this case, the maximum of $C_{v}$ increases with the number of IF neurons, $N$, faster than in the critical case, i.e. $\textrm{max}[C_{v}]\propto N^{a}$ with $a=1.70\pm 0.06$. Unlike $C_{v}$, the maximum of $\chi$ follows a scaling consistent with the critical case, i.e. $\textrm{max}[\chi]\propto N^{b}$ with $b=1.77\pm 0.12$.
@@
To conclude the analysis of the thermodynamics of the inferred Ising models, we examine the low temperature behavior of $C_{v}$ and $\chi$. For $T<T^{*}<1$, the numerical values of $C_{v}$ and $\chi$ depend on the initial spin configuration (see SI, Figs. S13 and S14). This interval $T<T^{*}$ is indicated by a gray-shaded background in Figs. 5 and 6 for the largest system sizes considered ($N=100$ neurons in the IF model and $N=60$ electrodes for experimental data). This behavior suggests the presence of a complex phase below the phase transition $T<T_{\textrm{max}}$, analogous to the spin-glass phase in the Sherrington-Kirkpatrick model [sherrington_solvable_1975], where many frustrated spin configurations are present and thermal fluctuations are insufficient to drive the system from a random initial state to the ground state within feasible Monte Carlo simulation times [newman_monte_1999, olsen_quality_2024]. The divergence of $C_{v}$ and $\chi$ as $T\to 0$ is therefore a numerical artifact, likely the result of averaging over different metastable states with varying energies $H(\bm{\sigma})$ and magnetizations $M(\bm{\sigma})$, resulting in large values for the fluctuations in Eqs. (20) and (21). The temperature $T^{*}$ depends on the system size, increasing with $N$, and on the state of the dynamics of the IF networks, being smallest for subcritical IF networks and largest for supercritical ones (see SI, Fig. S13 and Table S4). For low temperature $T\rightarrow 0$, Ising-like models with disorder in the parameters, such as the Sherrington–Kirkpatrick spin glass model, exhibit metastable states separated by free-energy barriers that scale with $N$ [aspelmeier_free-energy_2006]. If K‐pairwise models fitted to neuronal data would also exhibit similar free‐energy barriers that scale with $N$ at low temperature, then the shift of $T^{*}$ to higher temperatures with system size could be explained by the increasing thermal energy needed to overcome these barriers as $N$ increases.